OpenStudy (anonymous):

17-x over 4 = -10 may someone show me how to get the answer

OpenStudy (tyteen4a03):

\(17 - \frac{x}{4} = -10\) Multiply all terms by 4 to get rid of the denominator for the variable term.

OpenStudy (anonymous):

@tyteen4a03 i forgot to show what the problem looked like, sorry

OpenStudy (tyteen4a03):

In this case, multiply both sides by 4.

OpenStudy (phi):

if you start with \[ \frac{17-x}{4}= -10\] multiply both sides by 4 (this will simplify it a bit) \[\frac{\cancel{4}}{1}\cdot \frac{17-x}{\cancel{4}}= -10\cdot 4 \] \[ 17-x = -40 \] can you finish? One way to continue is add +x to both sides \[ 17 -x +x = -40 +x\] \[ 17 = -40 +x\] what is the final step?